Experimental Data is listed below for the reaction: A → 2B
Time (s) [B] (mol/L)
0.000 0.000
10.0 0.326
20.0 0.572
30.00 0.750
40.00 0.890
a) Prepare a graph from these data; connect the points with a smooth line; calculate the rate of
change of [B] for each 10 second interval fro 0 to 40.0 seconds. Does the rate of change
decrease from one interval to the next? Suggest a reason for the result.
b) How is the rate of change of [A] related to the rate of change of [B] in each time interval?
Calculate the rate of change of [A] fr the time interval from 10.0 to 20.0 seconds.
c) What is the instantaneous rate, ∆[B]/ ∆t when [B] = 0.750 mol/L?
1
Expert's answer
2020-11-14T13:52:03-0500

a) Yes, rate of change decreases from one interval to the next. This is because rate decreases as time increases

b) R = -d[A]/dt = d[B]/ 2 dt

From 10 to 20, Rate = 0.572-0.326 / 2(20-10) = 0.246/20= 0.0123

c) Rate = (0.750)/(2×30) = 0.750/60 = 0.0125